# Self-Reproduction and Evolution in Cellular Automata: 25 Years after Evoloops

### v1

@@ Abstract
The year of 2024 marks the 25th anniversary of the publication of evoloops, an evolutionary variant of Chris Langton’s self-reproducing loops which proved that Darwinian evolution of self-reproducing organisms by variation and natural selection is possible within deterministic cellular automata. Over the last few decades, this line of Artificial Life research has since undergone several important developments. Although it experienced a relative dormancy of activities for a while, the recent rise of interest in open-ended evolution and the success of continuous cellular automata models have brought researchers’ attention back to how to make spatio-temporal patterns self-reproduce and evolve within spatially distributed computational media. This article provides a review of the relevant literature on this topic over the past 25 years and highlights the major accomplishments made so far, the challenges being faced, and promising future research directions.
@@ 1 Introduction
The year of 2024 marks the 40th anniversary of Langton’s self-replicating loops and the 25th anniversary of self-reproducing evoloops. This line of research has remained, however, somewhat unpopular and unexplored ever since, likely because of the lack of rigorous theories, generalizability of models, and immediate applications to practical problem solving. Nonetheless, several important developments have since been made and, more recently, artificial life researchers have begun to pay attention again to how to make patterns self-replicate and evolve within spatially distributed computational media, potentially leading to open-ended evolution.
@@ 2 Background, Key Concepts and Advances
We first review Moore’s criterion for self-replication, distinguish trivial and non-trivial self-replication, draw the distinction between self-replication and self-reproduction, and then discuss the related concept of ‘individuals’ which these definitions tacitly presuppose and which will be important in relating self-reproduction to evolution. We then introduce von Neumann’s problems regarding self-reproduction and complexity, and his solutions or partial solutions to them. Langton’s views are then described and then von Neumann’s. This leads us to a brief review of evolution as it has been achieved in cellular automata.
@@ 2.1 Moore’s Criterion for Self-Replication and Beyond
E.F. Moore’s Criterion (1962) [32] uses the notion of a configuration surrounded by empty space in a cellular automaton as proxy for an ‘individual’. A configuration is a pattern of state activation within a bounded region of the space.
@@
Lones and Reggia [27] extended and made this criterion more elaborate; and then, using it to guide an evolutionary computation, they evolved self-replicating cellular automata configurations and their local update rules (unexpectedly self-replicating structures were moving!). Nehaniv and Dautenhahn [35] discuss further extension of Moore’s criterion so as to apply to self-reproduction, based on quantitative degree of spatiotemporial matching of configurations that may be non-exact copies.
2.2 Trivial and Non-Trivial Self-Replication
There are many cases though that satisfy Moore’s criteria but are borderline cases, not representative of living things, just patterns growing regularly, without discernable ‘individuals’. Their dynamics ensures replication without bound.
Examples include: (1) the growth of crystals or snowflakes, (2) spreading activation [^1] For example, a binary cellular automata with local states 0 ‘Off’ (quiescent), where a grid cell is unchanged in the next time step unless it has any neighbor that is 1 ‘On’ in which case its state is ‘On’ in the next time step., and (3) the modulo prime cellular automata of [3] which allow any 1D or 2D compact configuration to replicate using a special rule, where the new state of a cell is the modulo $p$ sum of states of neighbors where $p$ is any prime number and the local states are $\{0,1,\ldots,p-1\}$. This results in the reliable replication of any pattern, but this replication is a property of the ‘physics’ of the cellular space rather than dependent on any recognizable activity of individuals or their constitution.
A borderline trivial example of self-replication is the example of prions which are proteins taking conformations (shapes), either normal or pathological. A protein with the same amino acid sequence can fold up in different configurations to produce a typical ‘healthy’ conformation or one which causes disease and death. Pathological conformation is transmitted to healthy proteins by interaction with the proteins having the pathological, e.g. Creutzfeldt-Jakob disease (vCJD) or ‘mad cow disease’ in humans causing degeneration of the nervous system [13]. If prion proteins were to be considered ‘individuals’, that would be an argument for the single cells of a 1D cellular automata with spreading activation being individuals as well though they clearly should not be.
@@ 2.3 Self-Replication vs Self-Reproduction
We introduce definitions for self-replication and self-reproduction and their criteria in this section and then go on to describe how they arose and more recent developments.
@@ 2.3.1 Non-trivial Self-Replication and Self-Reproduction
Both von Neumann and Langton were concerned to exclude such trivial cases of self-replication and introduced different criteria for non-trivial self-reproduction (Sections 2.5 & 2.6). Self-replication in this review will refer to production of offspring that are exact copies of the parent. Langton created the first artificial non-trivial implemented self-replicating system in his 1984 paper [25]. Von Neumann gave the first rigorous descriptions and analyses of self-reproducing automata.
@@ 2.3.2 Individuals
The question of ‘What is an individual?’ is a deep and difficult one for biology and artificial life, that we must return to again and again. Individuality is intimately intertwined with self-reproduction and evolution. The origin, maintenance and evolution of individuality for multicellular entities with differentiated cells has been and continues to be an active and fruitful area of inquiry for biology [9, 31]. While Darwinian evolution describes dynamics of a population of individuals it does not define what is meant by ‘individual’. Indeed, Darwinian evolution works whether the ‘individuals’ are living organisms, viruses, programs in genetic programming, automata in evolutionary algorithms, or bitstrings in genetic algorithms. Moreover, the individuality of living organisms is not always discretely clear cut and often arises via the co-habitation of many ‘partners’, where the fertile concept of holobiont comprising a host organism and various microorganisms such as bacteria, viruses, and fungi that live within or the host, actively participating in its biology and sharing an intertwined fate with its lineage [28, 5].
@@ 2.4 Von Neumann’s Problem
The first one is solved explicitly in at least two fundamentally different ways by von Neumann in his book, most explicitly using cellular automata, while the second question is answered in principle by von Neumann’s “Fifth Lecture” [67] based on solution of the first question, but with no implementation being given or described in detail. Von Neuman’s conceptual solutions to the first part included both cellular automata and kinematic self-reproducer models, with the allegory of a robot on a ‘pond’ of components (warehouse) and is able to build a copy of itself by either of two different mechanisms, each in itself constituting a solution to the first part of von Neumann’s problem:
  1. 1.
    Self-examination (step-by-step build up copy by examining own structure). At no time does there exist a separate, complete description.
  2. 2.
    Build according to a description (program) from available components and copy program to ‘offspring’.
Articulation of these two mechanisms was remarkable at the time, with the second mechanism presaging the discovery of the role of DNA in life-as-we-know-it on earth: DNA as the molecular basis of heredity for life on earth was only discovered later in 1953 [70]. Indeed, DNA replication ‘blindly copies’ and protein biosynthesis ‘executes’ the code. So living things on earth satisfy Langton’s criterion for self-reproduction (see section 2.5 below).
McMullin [29] offers an insightful discussion of von Neumann’s pivotal contributions to understanding the evolutionary growth of complexity, the second part as outlined in von Neumann’s Fifth Lecture. Realization of von Neumann’s solution to the second problem had to wait for an implementation of these ideas in Sayama’s evoloops and Thomas S. Ray’s Tierra in the 1990s [44].
@@ 2.5 Christopher Langton and His Views
To avoid non-triviality in the characterization of self-reproduction, Langton [25] introduced criteria to exclude the trivial cases of self-replication described above (Section 2.2).
His implemented solution gives one answer to the first part of von Neumann’s problem, and using the second strategy for self-reproduction described by von Neumann, i.e., production of an offspring based on execution of instructions in a genome together with copying of that genome to the offspring.
To sum it up, his different criterion for self-reproduction is:
@@
Following Langton’s work, other researchers developed two directions: First, simplifying the replicators as much as possible: removal of one-side of the sheath around the circulating genome of the loop, and later complete removal of the sheath. While the steps in this simplification process preserve the dynamics of self-replication with the intention of meeting Langton’s criteria, it is hard to say that the resulting minimal replicators are really non-trivial when we compare them with the trivial replicators in Section 2.2. It seems that their complexity is in the eyes of the observer who knows the process that led to their invention. Second, adding complexity was pursued to give the self-replicating loops more computation power (stack automata capable of recognizing context free languages, or Turing complete computation), see the survey by Tempesti [64]. However, while the additional computation power allowed one to assert the replicators are computationally more complex than Langton’s loop, this complexity has not been used in any essential way in the self-replication process. It is an inessential “add-on” satisfying the mathematical criterion of higher complexity, but has not been used to achieve self-reproduction that supports inheritable variation nor has it served as a basis for evolution of further complexity.
@@ 2.6 Von Neumann’s Criterion for Self-Reproduction
In Theory of Self-Reproducing Automata [67, p.86], John von Neumann explicitly required a self-reproducer to have capacity for inheritable mutation. We can formulate:
  • Von Neumann’s Criterion for Self-Reproduction: A self-reproducer must have capacity for inheritable mutation in addition to the ability to replicate itself.
We use von Neumman’s definition and criterion for self-reproduction in this review. Langton’s loop is self-replicating and satisfies Langton’s criterion, but not von Neumann’s for a self-producing system. Moreover, von Neumann’s criterion is much more general in that it allows self-examination as an alternative mechanism for self-reproduction. This more general notion of self-reproduction is the key to solving the second part of von Neumann’s problem, as reviewed next.
@@ 3.1 Evoloops and Subsequent Studies through Mid-2010s
  • Finite lifetime - creating a turn-over of generations.
@@
Part of these aspects were successively addressed first by Langton’s self-reproducing loops [26] and then work of Sayama who solved the problem of freeing-up resources for generations using structurally dissolvale variants of Langton’s loop yielding a turn-over of generations [48]. Then, Sayama improved the robustness of state-transition rules to collisions and interactions between loops that made variability possible without catastrophic failure (inviability of loops). This resulted in evoloops [49, 50].
@@
The original evoloop [49, 50] was a 9-state 2D cellular automata model with von Neumann neighborhoods, derived from Langton’s self-reproducing loops [25]. Its state-transition function was revised thoroughly so that it would operate more robustly under a greater variety of local situations and that any undefined situation would generate a “dissolving” state that would propagate through contiguous active states and erase them from the space (structural dissolution). Direct interactions (collisions) of evoloops caused irregular situations during their replication process, naturally inducing variations of their genomes.
@@
Moreover, such variation in the genomes of the surviving loops lead to inheritability of this variation in a lineage. Differing reproductive success emerged intrinsically without any global fitness function in evoloops. The loops that were better and faster and able to survive and reproduce did so in the physics of the cellular automata universe (Figure 1) [49]. All state updates are local and the system includes no concept of “individual” or “selection” in its implementation. This was the first realization of a von Neumann self-reproducers evolving by natural selection. Remarkably, this was an instance of deterministic Darwinian evolution guaranteed to always play out the same way with the same initial conditions, with the trajectory of the evolving population depending on interactions between individuals for the generation of evolutionary novelty. The evoloop is also known to demonstrate substantial fault tolerance and, with slightly revised state transition function, abiogenesis from an initial configuration with no ancestor replicators [50].
@@
Figure 1: Spontaneous evolution of evoloops (image from [49]).
@@
The proposal of evoloops triggered several subsequent studies that implemented important developments. Nehaniv [36] expanded the model to asynchronously updated cellular automata and showed that evolution can occur even without synchronous updating. Sayama [52] introduced self-protection behaviors of individuals and showed that it would help promote the diversity of species. Salzberg and Sayama [46] conducted detailed genetic sequencing of all the individual self-replicators that appeared in simulations and found that their genotypic/phenotypic diversities were much greater than originally thought and they continued to evolve for a long period of time after the loops’ size reached the smallest level. Salzberg et al.  [45] also studied the evolutionary dynamics of evoloops in dynamic, hostile environments. Moreover, Oros and Nehaniv [39] presented a revised model called “sexyloop” in which loops engaged in sexual reproduction by exchanging genetic information when colliding with each other, and in later variant also showed that the capability of sexual reproduction can be evolutionarily maintained and impacts species diversity [40].
@@
A related approach studied at about the same time was to achieve evolution of self-replicators via the shape-encoding mechanism. It was originally proposed by Morita and Imai [34] for self-replication of various shapes in reversible cellular automata, but later adopted to promote spontaneous evolution of patterns through their spatial interactions [51, 60]. A concise review of research on self-replication and evolution in cellular automata up to mid-2000s can be found in [47].
@@
Since then, this line of research experienced a relative dormancy of activities for about a decade, probably because the topics of interest in the artificial life community became diversified and shifted more toward evolutionary robotics, neuroevolution, swarm intelligence, agent interactions, and others. During this “lost” decade, there was not much progress made in self-replicating and evolving cellular automata research, but some exceptions exist: In 2011, Yinusa and Nehaniv [71] showed explicitly that mutations to the type of von Neumann’s self-reproducing automata can result in non-viable offspring or failure to complete the reproduction process (lethal deleterious mutations), as well as heritabilty of neutral mutation, and successful complexity increase in lineage. However, the modifications to the inheritable information on the tape were ‘genetically engineered’ and not the result of a intrinsic process of variability generation. Another example is that, in 2013, Huang et al. [20] proposed self-reproducing loops that adapt their shapes to local spatial constraints.

### v2

@@ Abstract
The year of 2024 marks the 25th anniversary of the publication of evoloops, an evolutionary variant of Chris Langton’s self-reproducing loops which proved constructively that Darwinian evolution of self-reproducing organisms by variation and natural selection is possible within deterministic cellular automata. Over the last few decades, this line of Artificial Life research has since undergone several important developments. Although it experienced a relative dormancy of activities for a while, the recent rise of interest in open-ended evolution and the success of continuous cellular automata models have brought researchers’ attention back to how to make spatio-temporal patterns self-reproduce and evolve within spatially distributed computational media. This article provides a review of the relevant literature on this topic over the past 25 years and highlights the major accomplishments made so far, the challenges being faced, and promising future research directions.
@@ 1 Introduction
The year of 2024 marks the 40th anniversary of Langton’s self-replicating loops and the 25th anniversary of self-reproducing evoloops. This line of research has remained, however, somewhat unpopular and unexplored ever since, likely because of the lack of rigorous theories, generalizability of models, and immediate applications to practical problem solving. Nonetheless, several important developments have since been made and, more recently, artificial life researchers have begun to pay attention again to how to make patterns self-replicate and evolve within spatially distributed computational media (as reviewed in this article), potentially leading to open-ended evolution that keeps generating new forms of life indefinitely while exhibiting behavior of increasing complexity without converging toward some optimum [46, 47, 65, 69].
@@ 2 Background, Key Concepts and Advances
We mentioned above and will later return to the central themes of von Neumann’s seminal work on self-reproduction in automata (1948) [73]. We first highlight the various contributions of Moore, Langton, Sayama, and others, and their impact along the way in this narrative. For a number of years, work in this field focused somewhat over-restrictively on aspects emphasized in Langton’s work but now returns to the more general framing and original insights of von Neumann which embrace a broader notion of self-reproduction centered on the capacity for inheritable mutation in the production of offspring. First we review Moore’s criterion for self-replication, distinguish trivial and non-trivial self-replication, draw the distinction between self-replication and self-reproduction, and then discuss the related concept of ‘individuals’ which these definitions tacitly presuppose and which will be important in relating self-reproduction to evolution. [^1] The term individual occurs in three intertwined ways in this subject: first, there is the debate in biology of what constitutes an individual as a living self-maintaining organism (see Sec. 2.3.2); second, there is the notion of individual as an entity in a population undergoing a process of Darwinian evolution, which, like the primitive terms ‘line’ and ‘point’ in axiomatic geometry, may be treated as an undefined primitive entity in any system satisfying Darwin’s axioms for an evolutionary process (see Sec. 3.1); and finally there is the useful proxy notion of an individual in constructive artificial life studies as a configuration in cellular space surrounded by quiescent cells that replicates or reproduces, i.e., its instantiation at a single point in space-time (see Sec. 2.1). An underlying subtext of this review is to illustrate how the study of self-reproduction and evolution comes to integrate these notions of individual into one. We then introduce von Neumann’s problems regarding self-reproduction and complexity, and his solutions or partial solutions to them. Then Langton’s views are described and then von Neumann’s. This leads us to a brief review of evolution as it has been achieved in cellular automata in Section 3 and to future challenges and open problems in Section 4.
@@ 2.1 Moore’s Criterion for Self-Replication and Beyond
To support the study of self-reproduction from a logical perspective, E. F. Moore’s Criterion (1962) [36] uses the notion of a configuration surrounded by empty space in a cellular automaton as proxy for an ‘individual’. A configuration is a pattern of state activation within a bounded region of the space.
@@
Lones and Reggia [31] extended and made this criterion more elaborate; and then, using it to guide an evolutionary algorithm, they evolved self-replicating cellular automata configurations and their local update rules (unexpectedly self-replicating structures were moving!). Nehaniv and Dautenhahn [39] discuss further extension of Moore’s criterion so as to apply to self-reproduction, based on quantitative degree of spatiotemporial matching of configurations that may be non-exact copies.
2.2 Non-Triviality of Self-Replication
There are many cases though that satisfy Moore’s criteria or its physical analogues, but are borderline cases, not representative of living things, just patterns growing regularly, without discernable ‘individuals’. Their dynamics ensures replication without bound.
Examples include: (1) the growth of crystals or snowflakes, (2) spreading activation [^2] For example, a binary cellular automata with local states 0 ‘Off’ (quiescent), where a grid cell is unchanged in the next time step unless it has any neighbor that is 1 ‘On’ in which case its state is ‘On’ in the next time step. Although Moore’s criterion is technically violated here due to lack of surrounding quiescent cells, similar examples of single active cells for many elementary 1D cellular automata rules do satisfy the criterion.) , and (3) the modulo prime cellular automata [3] which allow any 1D or 2D compact configuration to replicate using a special rule, where the new state of a cell is the modulo $p$ sum of states of neighbors where $p$ is any prime number and the local states are $\{0,1,\ldots,p-1\}$. This results in the reliable replication of any pattern, but this replication is a property of the ‘physics’ of the cellular space rather than dependent on any recognizable activity of individuals or their constitution.
A borderline trivial example of self-replication is the example of prions which are proteins taking conformations (shapes), either normal or pathological. A protein with the same amino acid sequence can fold up in different configurations to produce a typical ‘healthy’ conformation or one which causes disease and death. Pathological conformation is transmitted to healthy proteins by interaction with the proteins having the pathological conformation, e.g. Creutzfeldt-Jakob disease (vCJD) or ‘mad cow disease’ in humans causing degeneration of the nervous system [14]. If prion proteins were to be considered ‘individuals’, that would be an argument for the single cells of a 1D cellular automaton with spreading activation being ‘individuals’ as well, though they clearly should not be.
@@ 2.3 Self-Replication vs Self-Reproduction
We have seen Moore’s criterion for self-replication and mentioned generalizations of his criterion to non-exact copies above, and we discussed some degenerate examples that do not reflect what one might expect for non-trivial self-replication or self-reproduction. However, the astute reader will observe that we still lack rigorous definitions distinguishing self-replication and self-reproduction. Therefore, we next articulate a definition for self-replication here and the two main competing criteria for self-reproduction due to Langton and von Neumann, respectively, in Sections 2.5 and 2.6, and then go on to describe how they arose and more recent developments in the sequel. It will turn out (1) that Langton’s criterion accepts some self-replicating systems, such as the Langton loop [27] or Sayama’s SDSR loop [54] (see below), although these are excluded by von Neumann’s criterion for self-reproduction as it requires viable inheritable variation, and (2) that Langton’s criterion is not broad enough to recognize self-reproduction in systems using self-examination to reproduce (a mechanism proposed by von Neumann – see Sec. 2.4). However, both Langton’s and von Neumann’s criteria do recognize evoloop [55], described below, as self-reproducing.
@@ 2.3.1 Non-trivial Self-Replication and Self-Reproduction
Both von Neumann and Langton were concerned to exclude such trivial cases of self-replication and introduced different criteria for non-trivial self-reproduction (Sections 2.5 & 2.6). Self-replication in this review will refer to production of offspring that are exact copies of the parent. This is consistent with Moore’s criterion for self-replication. Langton created the first artificial non-trivial implemented self-replicating system in his 1984 paper [27]. Von Neumann gave the first rigorous descriptions and analyses of self-reproducing automata [74].
@@ 2.3.2 Individuals
The question of ‘What is an individual?’ is a deep and difficult one for biology and artificial life, that we must return to again and again. Individuality is intimately intertwined with self-reproduction and evolution. The origin, maintenance and evolution of individuality for multicellular entities with differentiated cells has been and continues to be an active and fruitful area of inquiry for biology [10, 35]. While Darwinian evolution describes dynamics of a population of individuals, it does not define what is meant by ‘individual’. Indeed, Darwinian evolution works whether the ‘individuals’ are living organisms, viruses, programs in genetic programming, automata in evolutionary algorithms, or bitstrings in genetic algorithms. Moreover, the individuality of living organisms is not always discretely clear cut and often arises via the co-habitation of many ‘partners’; where we have the fertile concept of a holobiont comprising a host organism and various microorganisms such as bacteria, viruses, and fungi that live within the host or just in a close association (of different individuals – without any one identifiable as ‘host’), actively participating in its biology and sharing an intertwined fate with its lineage [32, 5].
@@ 2.4 Von Neumann’s Problem
The first one (vN-I) is solved explicitly in at least two fundamentally different ways by von Neumann in his book, most explicitly using cellular automata, while the second question (vN-II) is answered in principle by von Neumann’s “Fifth Lecture” [73] based on solution of the first question (vN-I), but with no implementation solving (vN-II) being given or described in detail. Von Neuman’s conceptual solutions to the first part included both cellular automata and kinematic self-reproducer models, with the allegory of a robot on a ‘pond’ of components (warehouse) that is able to build a copy of itself by either of two different mechanisms, each in itself constituting a solution to the first part of von Neumann’s problem (vN-I):
  1. Mech. 1.
    Self-examination: step-by-step build up of a copy (‘offspring’) by examining own structure. During this construction, at no time does there exist a separate complete description of the entity. [^3] One may consider this mechanism as reproduction in the absence of any separately identifiable component comprising a ‘genome’..
  2. Mech. 2.
    Build according to a description (program or ‘genome’) from available components and copy the program to the ‘offspring’.
Articulation of these two mechanisms was remarkable at the time, with the second one, Mech. 2, presaging the discovery of the role of DNA in life-as-we-know-it on earth: DNA structure as the molecular basis of heredity for life on earth was only discovered later in 1953 [76]. Indeed, DNA replication ‘blindly copies’ and protein biosynthesis ‘executes’ the code. So living things on earth satisfy Langton’s criterion for self-reproduction (see Section 2.5 below) and also von Neumann’s (Section 2.6), employing Mech. 2. Alternatively, self-reproduction via self-examination (Mech. 1) could be achieved, e.g., by an entity doing by a raster scan in 2D or 3D of itself and progressively printing a copy of itself as the scanning progresses, [^4] Possibly this could be done at a molecular level in physical analogues of self-reproduction with 3D scanning and printing. or by template-matching that copies the entire self, perhaps in two stages: producing a complementary strand and one complementary to the complement. Arguably, self-reproducing strands of RNA in a hypothesized pre-biotic RNA world are of this type [42].
McMullin [33] offers an insightful discussion of von Neumann’s pivotal contributions to understanding the evolutionary growth of complexity, the second part of von Neumann’s problem (vN-II), as outlined in von Neumann’s Fifth Lecture. Realization of von Neumann’s solution, i.e., harnessing Darwinian evolution of self-reproducers, to this second problem in artificial systems had to wait for novel mechanisms supporting inhertiable variation in H. Sayama’s evoloops [55] and T. S. Ray’s Tierra in the 1990s [49].
@@ 2.5 Christopher Langton and His Views
To ensure non-triviality in the characterization of self-reproduction, Langton [27] introduced the criterion presented below to exclude the trivial cases of self-replication described above (Section 2.2).
His implemented solution in the Langton loop [27] gives one answer to the first part of von Neumann’s problem (vN-I), using the second strategy for self-reproduction described by von Neumann (Mech. 2), i.e., production of an offspring based on execution of instructions in a genome together with copying of that genome to the offspring.
To sum it up, his specific criterion for self-reproduction is:
@@
Following Langton’s work, other researchers developed two directions: First, simplifying the replicators as much as possible: removal of one-side of the sheath around the circulating genome of the loop, and later complete removal of the sheath [8, 50]. While the steps in this simplification process preserve the dynamics of self-replication with the intention of meeting Langton’s criteria, it is hard to say that the resulting minimal replicators are really non-trivial when we compare them with the trivial replicators in Section 2.2. It seems that their complexity is in the eyes of the observer who knows the process that led to their invention. Second, adding complexity was pursued to give the self-replicating loops more computation power (stack automata capable of recognizing context free languages, or Turing complete computation); see the survey by Tempesti [70]. However, while the additional computation power allowed one to assert the replicators are computationally more complex than Langton’s loop, this complexity has not been used in any essential way in the self-replication process. It is an inessential “add-on” satisfying the mathematical criterion of higher complexity, but has not been used to achieve self-reproduction that supports inheritable variation nor has it served as a basis for evolution of further complexity.
While Langton’s criterion nicely captures a prerequisite to one of the methods (Mech. 2) of solving von Neumann’s problem (vN-I), it is silent about ensuring that complexity increase and an evolutionary process could be supported. That is, as much of the development following its introduction has shown, Langton’s criterion is neither necessary nor sufficient for addressing the second aspect of von Neumann’s problem (vN-II).
@@ 2.6 Von Neumann’s Criterion for Self-Reproduction
In Theory of Self-Reproducing Automata [73, p. 86], John von Neumann explicitly required a self-reproducer to have capacity for inheritable mutation. We can formulate:
  • Von Neumann’s Criterion for Non-trivial Self-Reproduction: A self-reproducer must have capacity for inheritable mutation in addition to the ability to replicate itself.
We use von Neumman’s definition and criterion for self-reproduction in this review. Langton’s loop is self-replicating and satisfies Langton’s criterion, but not von Neumann’s for a self-reproducing system. Moreover, von Neumann’s criterion is much more general in that it allows self-examination as an alternative mechanism for self-reproduction (Mech. 2), also discussed by Laing [29].
Does von Neumann’s criterion necessarily guarantee non-triviality? It does guarantee that the reproducer could in principle be an individual in an evolutionary process. The fact that some populations undergoes Darwinian evolution, say in a genetic algorithm or two-allele populations genetics model, does not necessarily entail much evolvability nor open-endedness. Similarly, von Neumann’s criterion might not necessarily lead to the evolution of complexity, nor even to any evolution at all – the context for Darwinian evolution may simply be lacking, as it is with von Neumann’s own example [73], where there are no sources of variability (see [73] for full details of von Neumann’s universal constructor self-reproducer and [77] for the variability and evolutionary considerations). However, von Neumann’s criterion is a necessary requisite for the (Darwinian) evolution of complexity in any population of self-reproducers, and thus a good criterion for non-triviality. This more general notion of self-reproduction is the key to solving the second part of von Neumann’s problem (vN-II) in full generality, as reviewed next.
@@ 3.1 Evoloops and Subsequent Studies through Mid-2010s
  • Finite Lifetime - a requisite for turn-over of generations.
@@
In cellular automata, finite space and resources arise from the use of bounded computational resources of the cellular automata space. [^5] This is automatic for bounded grids with finite update times (i.e, larger some positive lower-bound duration $\delta$) including those with periodic boundary conditions (“wrap-around”), and in fact for all present day real-world computational devices, even though they might be set up to emulate an unbounded environment. Other of these aspects (Darwinian axioms) were successively addressed first by Langton’s self-reproducing loops [28] (yielding Heredity, using Mech. 2, and Finite Lifetime) and then work of Sayama who solved the problem of freeing-up resources for generations using structurally dissolvable variants of Langton’s loop yielding a turn-over of generations [54]. Then, Sayama improved the robustness of state-transition rules to collisions and interactions between loops that made Variability possible without catastrophic failure (inviability of loops) despite some changes to the heritable information within the loops [55, 56]. This combination yielded Differing Fitness, i.e., reproductive success depending in part on inherited characteristics and resulted in evoloops.
@@
The original evoloop system [55, 56] was a 9-state 2D cellular automata model with von Neumann (5-cell) neighborhoods [74], [^6] The local synchronous update of the state at each location (cell) depends on the state of the cells in the 2D grid located above, below, left, and right of a given cell, and on the state of the cell itself. derived from Langton’s self-reproducing loops [27]. Its state-transition function was revised thoroughly so that it would operate more robustly under a greater variety of local situations [55] and that any undefined situation would generate a “dissolving” state that would propagate through contiguous active states and erase them from the space (structural dissolution) [54, 55]. Turn-over of generations is made possible in this way by freeing up space which otherwise would become clogged with the debris and remnants from no longer active individuals (as in loop models prior to Sayama’s), and the robustness frequently makes inheritable variation possible in loops still able to reproduce.
@@
In the evoloop model, direct interactions (collisions) of loops caused irregular situations during their replication process, naturally inducing variations of their genomes. Such variations in the genomes of the surviving loops lead to inheritability of this variation in a lineage. Differing reproductive success emerged intrinsically without any global fitness function in evoloops. The loops that were better and faster and able to survive and reproduce did so in the physics of the cellular automata universe (Figure 1) [55]. All state updates are local and the system includes no concept of “individual” or “selection” in its implementation. This was the first realization of von Neumann self-reproducers evolving by natural selection. Remarkably, this was an instance of determinisic Darwinian evolution guaranteed to always play out the same way with the same initial conditions, with the trajectory of the evolving population depending on interactions between individuals for the generation of evolutionary novelty. The evoloop is also known to demonstrate substantial fault tolerance and, with slightly revised state transition function, abiogenesis from an initial configuration with no ancestor replicators [56].
@@
Figure 1: Spontaneous evolution of evoloops (image from [55]). Time indicates the number of updates applied.
@@
The advent of evoloops triggered several subsequent studies that implemented important developments. Nehaniv [40] expanded the model to asynchronously updated cellular automata and showed that evolution can occur even without synchronous updating. Sayama [58] introduced self-protection behaviors of individuals and showed that it would help promote the diversity of species. Salzberg and Sayama [52] conducted detailed genetic sequencing of all the individual self-replicators that appeared in simulations and found that their genotypic/phenotypic diversities were much greater than originally thought and they continued to evolve for a long period of time after the loops’ size reached the smallest level. Salzberg et al.  [51] also studied the evolutionary dynamics of evoloops in dynamic, hostile environments. A concise review of research on self-replication and evolution in cellular automata up to mid-2000s can be found in [53]. Moreover, Oros and Nehaniv [44] then presented a revised model called “sexyloop” in 2007 in which loops engaged in sexual reproduction by exchanging genetic information when colliding with each other; and in a later variant in 2009, where the capacity for sex was made inheritable (rather than instrinsic in the physics of the update rules), they also showed that the capability of sexual reproduction can be evolutionarily maintained and impacts species diversity [45].
@@
A related approach studied at about the same time was to achieve evolution of self-replicators via the shape-encoding mechanism. It was originally proposed by Morita and Imai [38] for self-replication of various shapes in reversible cellular automata, but later adopted to promote spontaneous evolution of patterns through their spatial interactions [57, 66].
@@
Since then, this line of research experienced a relative dormancy of activities for about a decade, probably because the topics of interest in the artificial life community became diversified and shifted more toward evolutionary robotics, neuroevolution, swarm intelligence, agent interactions, and others. During this “lost” decade, there was not much progress made in self-reproducing and evolving cellular automata research, but some exceptions exist: In 2011, Yinusa and Nehaniv [77] showed explicitly that mutations to the tape (genetic information encoding its construction) of von Neumann’s self-reproducing automata can result in non-viable offspring or failure to complete the reproduction process (lethal deleterious mutations), as well as heritabilty of neutral mutation, and also in successful complexity increase in a lineage. However, the modifications to the inheritable information on the tape were ‘genetically engineered’ and not the result of a intrinsic process of variability generation. Another example is that, in 2013, Huang et al. [21] proposed self-reproducing loops that adapt their shapes to local spatial constraints.